OpenStudy (anonymous):

find the value of x. the exterior angles are 2x, 3x, (x+10), (4x-15), and (2x+20) degrees

OpenStudy (anonymous):

Assuming we're talking about the exterior angles of a polygon, the sum of all exterior angles is 360 degrees, so... 2x + 3x + (x + 10) + (4x - 15) + (2x + 20) = 360 12x + 15 = 360 12x = 345 x = 345/12 = 115/4 = 28.75

OpenStudy (anonymous):

thank you